# LtU-ILI: An All-in-One Framework for Implicit Inference in Astrophysics and Cosmology

### v1

@@ 4.1 Toy Problem
where $k_{i}=(2i/3)-3$ and $i\in[0,9]$. Here, the data vector $\bm{x}:=\{x_{i}\}_{i=0}^{9}\in\mathbb{R}^{10}$ is composed of a sinusoidal signal with phase $\phi_{0}$, a quadratic signal of amplitude $\phi_{1}$, and an i.i.d. noise component $\epsilon_{i}\sim\mathcal{N}(0,1)$. We further decompose ${\bm{\phi}}$ into our target parameters $\bm{\theta}\in\mathbb{R}^{3}$ via $\phi_{0}=\theta_{0}+\theta_{1}$ and $\phi_{1}=\theta_{1}-3\theta_{2}^{2}$. We expect our inference engine to forward constraints on latent variables $\bm{\phi}$ onto target variables $\bm{\theta}$. We assume a standard normal prior on each of our parameters of interest, $p(\theta_{i})=\mathcal{N}(0,1)$. This is a challenging inference task because the likelihood is highly non-linear and the posteriors exhibit strong degeneracies. However, learning parameters from such complex, correlated data vectors is a common problem in astronomy and cosmology (e.g., Dax et al.(2021a)Dax, Green, Gair, Macke, Buonanno, and Schölkopf; Modi et al.(2023)Modi, Pandey, Ho, Hahn, Blancard, and Wandelt; Bartlett et al.(2023)Bartlett, Kammerer, Kronberger, Desmond, Ferreira, Wandelt, Burlacu, Alonso, and Zennaro).
@@ 4.2 Benchmarking
Figure 4: Prediction error as a function of simulation budget for various LtU-ILI training methodologies on the benchmark SLCP inference problem (Papamakarios et al.(2019)Papamakarios, Sterratt, and Murray). For clarity, plotted points for different models at the same simulation budget are slightly offset horizontally. The Rejection-ABC method is shown on all subplots as a baseline. Error is defined in terms of the Classifier 2-sample Test (C2ST; Lopez-Paz and Oquab(2016)) metric. A lower C2ST indicates a more accurate inferred posterior, with 0.5 being the optimal score. C2ST values are shown at their median and central 95% confidence interval, calculated over ten independent runs.
We benchmark all models in LtU-ILI on the standard “Simple Likelihood Complex Posterior” problem (SLCP; Papamakarios et al.(2019)Papamakarios, Sterratt, and Murray) common in simulation-based inference experiments (e.g., Greenberg et al.(2019)Greenberg, Nonnenmacher, and Macke; Lueckmann et al.(2021)Lueckmann, Boelts, Greenberg, Goncalves, and Macke; Ramesh et al.(2022)Ramesh, Lueckmann, Boelts, Tejero-Cantero, Greenberg, Gonçalves, and Macke). SLCP has a fairly straightforward, unimodal likelihood but a very complex, multimodal posterior distribution, making it a difficult, but attainable benchmark for implicit inference. We follow the implementation of SLCP in Lueckmann et al.(2021)Lueckmann, Boelts, Greenberg, Goncalves, and Macke, wherein we train various NPE/NLE/NRE methods and compare their resultant posteriors to reference posteriors determined with long-run HMC chains. We implement this for all available NPE/NLE/NRE methods and their sequential analogs in each of the pydelfi, sbi, and lampe backends. We also include benchmarking of Rejection ABC as a baseline. Throughout all backends and engines, we use the same architectures as in Lueckmann et al.(2021)Lueckmann, Boelts, Greenberg, Goncalves, and Macke, which were determined through exhaustive hyperparameter searches.
@@
In Figure 4, we show the performance of each LtU-ILI engine on the SLCP benchmark as a function of the simulation budget used for generating the training set. As we would expect, the accuracy of inferred posteriors increases with the number of given simulations across all models. We observe that all models strongly outperform the Rejection ABC baseline due to their efficient use of simulation data. For the SLCP benchmark, we observe that NLE methods perform the best, NPE methods are a close second, and NRE methods have the worst performance. However, we note that the SLCP by definition has a simple, easy-to-learn likelihood, encouraging strong NLE performance. This hierarchy of performance could be entirely different for other problems. Sequential methods (SNPE/SNLE/SNRE) show very little improvement for small simulation budgets ($10^{3}$), but show massive gains for large budgets ($10^{5}$), especially for the NLE methods. This is likely because, for small training datasets, the posteriors are poorly constrained and barely deviate from the prior. However, for the tight posterior constraints in large budget runs, sequential methods are able to more easily focus simulation resources on small regions of parameter space. Next, we note that the lampe-NPE implementation largely mirrors that of sbi-NPE, and pydelfi-NLE slightly outperforms sbi-NLE for a medium simulation budget ($10^{4}$). Although hyperparameters (i.e., architectures, learning rate, validation schema) are fixed for these experiments, slight differences in backend implementations might lead to better or worse performance. Lastly, as a check, the performances of all sbi engines closely match those found in Lueckmann et al.(2021)Lueckmann, Boelts, Greenberg, Goncalves, and Macke, suggesting our implementation of SLCP is correct.
@@ 5 Science Experiments
In this section, we demonstrate various applications of LtU-ILI for common astrophysics and cosmology problems. These experiments are designed to display the diverse functionality of the code, including using CNN embedding networks (Section 5.1), applying posterior coverage validation (Section 5.2), incorporating graph neural network layers (Section 5.3), using multiround inference (Section 5.4), quantifying information in different data sources (Section 5.5), and integrating information-optimal pretrained embeddings (Section 5.6). We note that these experiments are intended to illuminate interesting pathways for the scientific application of LtU-ILI and not necessarily make significant scientific conclusions on their subject material. For readability, experiment descriptions are kept short, but the code and data are made publicly available through the codebase.
@@ 5.1 X-ray Mass Estimation of Galaxy Clusters
Figure 5: Inference of galaxy cluster $M_{500c}$ mass from eROSITA-like mock X-ray observations, as in Ho et al.(2023)Ho, Soltis, Farahi, Nagai, Evrard, and Ntampaka. The left four subplots show random examples of single-band X-ray observables in terms of their sky-projected surface brightness and target mass (in $h^{-1}M_{\odot}$). The right plot is the true vs. predicted log-mass estimates on the test set, where we show the median and central 68% confidence interval of our neural posteriors. For readability, test points in the right plot have been randomly subsampled.
@@
A prime functionality of LtU-ILI is its flexibility to adapt to many modalities of data. Through the use of a customizable embedding network, we process image data to estimate the mass of galaxy clusters from X-ray observations. Our inference problem mirrors that of Ho et al.(2023)Ho, Soltis, Farahi, Nagai, Evrard, and Ntampaka, wherein we try to estimate the $M_{\rm 500c}$ mass of a galaxy cluster from $128\times 128$ images of sky-projected X-ray photon counts. To train and test our model, we use a dataset of $3,285$ mock X-ray observations of clusters derived from the Magneticum hydrodynamical simulation (Dolag et al.(2016)Dolag, Komatsu, and Sunyaev) designed to mimic those of the eROSITA telescope (Soltis et al.(2022)Soltis, Ntampaka, Wu, ZuHone, Evrard, Farahi, Ho, and Nagai). These mock observations are single-band images of the X-ray photon emission from the hot intra-cluster medium (ICM) and include realistic simulation of the contaminating systematics which afflict real X-ray measurements, including cluster morphology, background emission, telescope response, and AGN sources. Examples of these observables are shown in Figure 5. A good estimator of cluster mass needs to be able to disentangle the source X-ray signal from the noise and peripheral emission, while also understanding the physical connection between X-ray emission, ICM gas content, and system mass.
@@
To recreate this example, we perform NPE using an embedding network of the same convolutional architecture as Ho et al.(2023)Ho, Soltis, Farahi, Nagai, Evrard, and Ntampaka. That is, instead of the last dense layer mapping to a single point estimate, we forward the final embedding into an NDE. We construct an ensemble of four models, each with the same embedding architecture, but two with Masked Autoregressive Flow (MAF Papamakarios et al.(2017)Papamakarios, Pavlakou, and Murray) NDEs and two with MDNs. We then train the models from scratch using the NDE losses, instead of the mean-squared-error (MSE) loss of Ho et al.(2023)Ho, Soltis, Farahi, Nagai, Evrard, and Ntampaka. As in the reference, we assume a uniform prior on cluster mass. We split our mock catalog into 90% training data and 10% test data, and present our performance results on the latter. The entire training and testing procedure takes about fifteen minutes on an Nvidia V100 GPU.
@@
The true vs. predicted mass estimates for this model are shown in Figure 5. Using the original architecture, we achieve a test scatter on the mean prediction of $0.0782\ {\rm dex}$. When compared to the $0.0773\ {\rm dex}$ scatter of the single-band X-ray images in Ho et al.(2023)Ho, Soltis, Farahi, Nagai, Evrard, and Ntampaka, we achieve a very similar level of information extraction, though our uncertainty is slightly higher than the original. We attribute this to the fact that model ensembling slightly inflates predictive uncertainty (Hermans et al.(2022)Hermans, Delaunoy, Rozet, Wehenkel, Begy, and Louppe). However, we remark that by slightly increasing the number of convolutional filters in the first two layers, we observed mean scatters as low as $0.071\ {\rm dex}$, suggesting that the NPEs are capable of state-of-the-art constraints on physical problems when given the right architecture. We also note that hyperparameter search through neural architectures is made efficient and accessible through LtU-ILI’s configuration-based interface.
@@ 5.2 Quijote Dark Matter Power Spectrum
Refer to caption (a) True vs. Predicted Refer to caption (b) Marginal Percentile Coverage Test
(c) TARP Multivariate Coverage Test
Figure 6: Inference of three cosmological parameters $\Omega_{m}$, $h$, and $\sigma_{8}$ from halo power spectrum multipoles in Quijote simulations using SBI. In (a) we show the comparison between true and predicted value of the parameters in the test simulations. Marginal error bars are shown at the central $68\%$ confidence intervals. In (b) and (c) we show the coverage tests using the P-P plots as well as using TARP method. These show bootstrapping error bars at $68\%$ and $95\%$, indicative of typical variations in an ideal estimator for a finite test set.
[…]
We train an ensemble of six networks, each using the MAF architecture with three transformations and 10 hidden features. We use 1800 simulations to train the network, retaining the remaining 200 to test its performance as shown in Figure 6. We then use VI sampling to obtain the posterior of the parameters from this trained likelihood ensemble, weighted by their validation losses.
@@ 5.3 Field-level Inference with Graph Neural Networks
We next demonstrate LtU-ILI’s ability to handle different data structures by also including a point cloud dataset example. In particular, we train a neural posterior estimator to recover the cosmological parameters $\Omega_{m}$ and $\sigma_{8}$ from the positions of the $10,000$ heaviest dark matter haloes in the Quijote simulations (Villaescusa-Navarro et al.(2020)Villaescusa-Navarro, Hahn, Massara, Banerjee, Delgado, Ramanah, Charnock, Giusarma, Li, Allys, Brochard, Uhlemann, Chiang, He, Pisani, Obuljen, Feng, Castorina, Contardo, Kreisch, Nicola, Alsing, Scoccimarro, Verde, Viel, Ho, Mallat, Wandelt, and Spergel). The high resolution information present in the discrete halo positions has been shown to constrain cosmological parameters to extremely high precision (e.g., Makinen et al.(2022b)Makinen, Charnock, Lemos, Porqueres, Heavens, and Wandelt; Cuesta-Lazaro and Mishra-Sharma(2023); de Santi et al.(2023b)de Santi, Shao, Villaescusa-Navarro, Abramo, Teyssier, Villanueva-Domingo, Ni, Anglés-Alcázar, Genel, Hernández-Martínez, Steinwandel, Lovell, Dolag, Castro, and Vogelsberger).
As an embedding network, we use a graph neural network. We construct a graph where dark matter haloes act as nodes connected to any other dark matter halo separated by a distance smaller than $20\,{\rm Mpc}/h$.
We train an ensemble of four networks, each one using a MAF architecture with five transformations and $50$ hidden features. The graph neural network is a message-passing neural network with node and edge neural networks being two multi-layer perceptrons with $3$ layers of $128$ hidden units each. The network takes as inputs 3D distances between nodes, and their corresponding modulus.
In Figure 7, we show an example point cloud together with the resulting graph neural network inference. The graph neural network obtains well-calibrated and informative posteriors, consistent with true values. We note that the constraints are not as tight as those produced for the Quijote dark matter power spectrum in Figure 6. This is likely because the graph only targets the most massive halos, i.e., a more physically realistic probe, while the results in Figure 6 access the full, high-resolution dark matter distribution.
@@ 5.4 Multi-round Inference for Gravitational Waves
Refer to caption (a) Example of GW signal in source and detector frame. Refer to caption (b) Multi-round training and validation posterior probability.
@@
Over recent years, several studies based on ILI methods have addressed the problem of reconstructing parameters of merger events from gravitational wave (GW) signals (Cutler and Flanagan(1994); Thrane and Talbot(2019)). Initial advancements were marked by significant contributions to posterior estimation using normalizing flows, as highlighted in Green et al.(2020)Green, Simpson, and Gair; Green and Gair(2020). Subsequent research (DINGO: Dax et al.(2021a)Dax, Green, Gair, Macke, Buonanno, and Schölkopf) introduced variations of NPE taking advantage of Group Equivariant Neural Posterior Estimation (GNPE: Dax et al.(2021b)Dax, Green, Gair, Deistler, Schölkopf, and Macke) or Flow Matching Posterior Estimation (Wildberger et al.(2023)Wildberger, Dax, Buchholz, Green, Macke, and Schölkopf). More recently, Crisostomi et al.(2023)Crisostomi, Dey, Barausse, and Trotta showcased a similar application of multi-round NPE to estimate ringdown parameters of GW150914. Delaunoy et al.(2020)Delaunoy, Wehenkel, Hinderer, Nissanke, Weniger, Williamson, and Louppe and Bhardwaj et al.(2023)Bhardwaj, Alvey, Miller, Nissanke, and Weniger also serve as recent illustrations of applying NRE to GW analyses – the latter incorporating multi-round inference as well.
@@
In this example, we explore the application of multi-round inference to constrain a reduced set of gravitational wave (GW) parameters. Our forward-modeled data vector is constructed by merging the simulated frequency-domain strain signals from each of the Hanford (H1) and Livingstone (L1) detectors. We employ the IMRPhenomPv2 algorithm (Khan et al.(2016)Khan, Husa, Hannam, Ohme, Pürrer, Forteza, and Bohé) with a frequency resolution of $\Delta f=0.125\textrm{ Hz}$, resulting in an input data vector of $\textrm{dim}(\bm{x})=2\times 7872$. An illustration of this data vector is presented in Figure 8a for clarity. We stress that, in contrast to likelihood-based sampling methods implemented in Bilby (Ashton et al.(2019)Ashton, Hübner, Lasky, Talbot, Ackley, Biscoveanu, Chu, Divakarla, Easter, Goncharov, Hernandez Vivanco, Harms, Lower, Meadors, Melchor, Payne, Pitkin, Powell, Sarin, Smith, and Thrane; Romero-Shaw et al.(2020)Romero-Shaw, Talbot, Biscoveanu, D’Emilio, Ashton, Berry, Coughlin, Galaudage, Hoy, Hübner, Phukon, Pitkin, Rizzo, Sarin, Smith, Stevenson, Vajpeyi, Arène, Athar, Banagiri, Bose, Carney, Chatziioannou, Clark, Colleoni, Cotesta, Edelman, Estellés, García-Quirós, Ghosh, Green, Haster, Husa, Keitel, Kim, Hernandez-Vivanco, Magaña Hernandez, Karathanasis, Lasky, De Lillo, Lower, Macleod, Mateu-Lucena, Miller, Millhouse, Morisaki, Oh, Ossokine, Payne, Powell, Pratten, Pürrer, Ramos-Buades, Raymond, Thrane, Veitch, Williams, Williams, and Xiao), which only necessitate the signal component of the forward model for the likelihood computation, our ILI approach requires the inclusion of noise realizations in the training data vector.
@@
We illustrate a simplified application of LtU-ILI for NPE on GW signals, fixing parameters related to the time and geometry of the detection and focusing on a set of four parameters: chirp mass ($\mathcal{M}$), mass ratio ($q$), effective aligned spin ($\chi_{\textrm{eff}}$) and luminosity distance ($d_{L}$). These parameters, chosen for their reduced correlation and significance in GW parameter estimation (e.g., Veitch et al.(2015)Veitch, Raymond, Farr, Farr, Graff, Vitale, Aylott, Blackburn, Christensen, Coughlin, Del Pozzo, Feroz, Gair, Haster, Kalogera, Littenberg, Mandel, O’Shaughnessy, Pitkin, Rodriguez, Röver, Sidery, Smith, Van Der Sluys, Vecchio, Vousden, and Wade; Vitale et al.(2017)Vitale, Lynch, Raymond, Sturani, Veitch, and Graff), are derived from the masses of the coalescing objects – with $\mathcal{M}\equiv(m_{1}m_{2})^{\frac{3}{5}}/(m_{1}+m_{2})^{\frac{1}{5}}$ and $q\equiv m_{2}/m_{1}\leq 1$ – and spin characteristics (Racine(2008)).
@@
We established baseline values for our simulated event $\bm{x_{0}}$ at {$d_{L}=390\ \textrm{Mpc},q=0.79,\mathcal{M}=30.2\ M_{\odot},\chi_{\textrm{eff}}=-0.09$)}, akin to the GW150914 event as detailed in the first column of Table 1 by Abbott et al.(2016)Abbott et al.. Following the procedure described in Karathanasis et al.(2023)Karathanasis, Revenu, Mukherjee, and Stachurski, the defined priors are uniform ranging from [100.0 $1.0100.0\textrm{ Mpc},0.2,12.0\,M_{\odot},-$ 1.0] to [1000.0 $1.01000.0\textrm{ Mpc},1.0,45.0\,M_{\odot},$ 1.0], with narrower margins for chirp mass and luminosity distance to facilitate a straightforward example. For density estimation, we utilize the sbi backend with a MAF and a 1D convolutional embedding network, undergoing five rounds of training with 1000 simulations per round.

### v2

@@ 4.1 Toy Problem
where $k_{i}=(2i/3)-3$ and $i\in[0,9]$. Here, the data vector $\bm{x}:=\{x_{i}\}_{i=0}^{9}\in\mathbb{R}^{10}$ is composed of a sinusoidal signal with phase $\phi_{0}$, a quadratic signal of amplitude $\phi_{1}$, and an i.i.d. noise component $\epsilon_{i}\sim\mathcal{N}(0,1)$. We further decompose ${\bm{\phi}}$ into our target parameters $\bm{\theta}\in\mathbb{R}^{3}$ via $\phi_{0}=\theta_{0}+\theta_{1}$ and $\phi_{1}=\theta_{1}-3\theta_{2}^{2}$. We expect our inference engine to forward constraints on latent variables $\bm{\phi}$ onto target variables $\bm{\theta}$. We assume a standard normal prior on each of our parameters of interest, $p(\theta_{i})=\mathcal{N}(0,1)$. This is a challenging inference task because the likelihood is highly non-linear and the posteriors exhibit strong degeneracies. However, learning parameters from such complex, correlated data vectors is a common problem in astronomy and cosmology (e.g., Dax et al., 2021a; Modi et al., 2023; Bartlett et al., 2023).
@@ 4.2 Benchmarking
Figure 4: Prediction error as a function of simulation budget for various LtU-ILI training methodologies on the benchmark SLCP inference problem (Papamakarios et al., 2019). For clarity, plotted points for different models at the same simulation budget are slightly offset horizontally. The Rejection-ABC method is shown on all subplots as a baseline. Error is defined in terms of the Classifier 2-sample Test (C2ST; Lopez-Paz and Oquab, 2016) metric. A lower C2ST indicates a more accurate inferred posterior, with 0.5 being the optimal score. C2ST values are shown at their median and central 95% confidence interval, calculated over ten independent runs.
We benchmark all models in LtU-ILI on the standard “Simple Likelihood Complex Posterior” problem (SLCP; Papamakarios et al., 2019) common in simulation-based inference experiments (e.g., Greenberg et al., 2019; Lueckmann et al., 2021; Ramesh et al., 2022). SLCP has a fairly straightforward, unimodal likelihood but a very complex, multimodal posterior distribution, making it a difficult, but attainable benchmark for implicit inference. We follow the implementation of SLCP in Lueckmann et al. (2021), wherein we train various NPE/NLE/NRE methods and compare their resultant posteriors to reference posteriors determined with long-run HMC chains. We implement this for all available NPE/NLE/NRE methods and their sequential analogs in each of the pydelfi, sbi, and lampe backends. We also include benchmarking of Rejection ABC as a baseline. Throughout all backends and engines, we use the same architectures as in Lueckmann et al. (2021), which were determined through exhaustive hyperparameter searches.
@@
In Figure 4, we show the performance of each LtU-ILI engine on the SLCP benchmark as a function of the simulation budget used for generating the training set. As we would expect, the accuracy of inferred posteriors increases with the number of given simulations across all models. We observe that all models strongly outperform the Rejection ABC baseline due to their efficient use of simulation data. For the SLCP benchmark, we observe that NLE methods perform the best, NPE methods are a close second, and NRE methods have the worst performance. However, we note that the SLCP by definition has a simple, easy-to-learn likelihood, encouraging strong NLE performance. This hierarchy of performance could be entirely different for other problems. Sequential methods (SNPE/SNLE/SNRE) show very little improvement for small simulation budgets ($10^{3}$), but show massive gains for large budgets ($10^{5}$), especially for the NLE methods. This is likely because, for small training datasets, the posteriors are poorly constrained and barely deviate from the prior. However, for the tight posterior constraints in large budget runs, sequential methods are able to more easily focus simulation resources on small regions of parameter space. Next, we note that the lampe-NPE implementation largely mirrors that of sbi-NPE, and pydelfi-NLE slightly outperforms sbi-NLE for a medium simulation budget ($10^{4}$). Although hyperparameters (i.e., architectures, learning rate, validation schema) are fixed for these experiments, slight differences in backend implementations might lead to better or worse performance. Lastly, as a check, the performances of all sbi engines closely match those found in Lueckmann et al. (2021), suggesting our implementation of SLCP is correct.
@@ 5 Science Experiments
In this section, we demonstrate various applications of LtU-ILI for common astrophysics and cosmology problems. These experiments are designed to display the diverse functionality of the code, including using CNN embedding networks (Section 5.1), applying posterior coverage validation (Section 5.2), incorporating graph neural network layers (Section 5.3), using multiround inference (Section 5.4), quantifying information in different data sources (Section 5.5), and integrating information-optimal embeddings (Section 5.6). We note that these experiments are intended to illuminate interesting pathways for the scientific application of LtU-ILI and not necessarily make significant scientific conclusions on their subject material. For readability, experiment descriptions are kept concise, but the code and data are made publicly available through the codebase.
@@ 5.1 X-ray Mass Estimation of Galaxy Clusters
Figure 5: Inference of galaxy cluster $M_{500c}$ mass from eROSITA-like mock X-ray observations, as in Ho et al. (2023). The left four subplots show random examples of single-band X-ray observables in terms of their sky-projected surface brightness and target mass (in $h^{-1}M_{\odot}$). The right plot is the true vs. predicted log-mass estimates on the test set, where we show the median and central 68% confidence interval of our neural posteriors. For readability, test points in the right plot have been randomly subsampled.
@@
A prime functionality of LtU-ILI is its flexibility to adapt to many modalities of data. Through the use of a customizable embedding network, we process image data to estimate the mass of galaxy clusters from X-ray observations. Our inference problem mirrors that of Ho et al. (2023), wherein we try to estimate the $M_{\rm 500c}$ mass of a galaxy cluster from $128\times 128$ images of sky-projected X-ray photon counts. To train and test our model, we use a dataset of $3,285$ mock X-ray observations of clusters derived from the Magneticum hydrodynamical simulation (Dolag et al., 2016) designed to mimic those of the eROSITA telescope (Soltis et al., 2022). These mock observations are single-band images of the X-ray photon emission from the hot intra-cluster medium (ICM) and include realistic simulation of the contaminating systematics which afflict real X-ray measurements, including cluster morphology, background emission, telescope response, and AGN sources. Examples of these observables are shown in Figure 5. A good estimator of cluster mass needs to be able to disentangle the source X-ray signal from the noise and peripheral emission, while also understanding the physical connection between X-ray emission, ICM gas content, and system mass.
@@
To recreate this example, we perform NPE using an embedding network of the same convolutional architecture as Ho et al. (2023). That is, instead of the last dense layer mapping to a single point estimate, we forward the final embedding into an NDE. We construct an ensemble of four models, each with the same embedding architecture, but two with Masked Autoregressive Flow (MAF; Papamakarios et al., 2017) NDEs and two with MDNs. We then train the models from scratch using the NDE losses, instead of the mean-squared-error (MSE) loss of Ho et al. (2023). As in the reference, we assume a uniform prior on cluster mass. We split our mock catalog into 90% training data and 10% test data, and present our performance results on the latter. The entire training and testing procedure takes about fifteen minutes on an Nvidia V100 GPU.
@@
The true vs. predicted mass estimates for this model are shown in Figure 5. Using the original architecture, we achieve a test scatter on the mean prediction of $0.0782\ {\rm dex}$. When compared to the $0.0773\ {\rm dex}$ scatter of the single-band X-ray images in Ho et al. (2023), we achieve a very similar level of information extraction, though our uncertainty is slightly higher than the original. We attribute this to the fact that model ensembling slightly inflates predictive uncertainty (Hermans et al., 2022). However, we remark that by slightly increasing the number of convolutional filters in the first two layers, we observed mean scatters as low as $0.071\ {\rm dex}$, suggesting that the NPEs are capable of state-of-the-art constraints on physical problems when given the right architecture. We also note that hyperparameter search through neural architectures is made efficient and accessible through LtU-ILI’s configuration-based interface.
(a) True vs. Predicted
(b) Marginal Percentile Coverage Test
(c) TARP Multivariate Coverage Test
Figure 6: Inference of three cosmological parameters $\Omega_{m}$, $h$, and $\sigma_{8}$ from halo power spectrum multipoles in Quijote simulations using LtU-ILI. In (a) we show the comparison between true and predicted value of the parameters in the test simulations. Marginal error bars are shown at the central $68\%$ confidence intervals. In (b) and (c) we show the coverage tests using the P-P plots as well as using TARP method. These show bootstrapping error bars at $68\%$ and $95\%$, indicative of typical variations in an ideal estimator for a finite test set.
@@ 5.2 Quijote Dark Matter Power Spectrum
Another prime utility of LtU-ILI is its capacity to seamlessly integrate ILI training procedures with sampling and validation metrics. A common inference benchmark in the field of cosmology is inferring cosmological parameters from the matter power spectrum (e.g., Cooray and Sheth, 2002; Spurio Mancini et al., 2022; Hahn et al., 2023b). In this problem, we run simulations of the matter evolution in the universe for various cosmological models, measure observational summary statistics such as the matter power spectrum, and then try to connect these observations back to constraints on cosmology. Here, we demonstrate this application by predicting the posterior distribution of three cosmological parameters using the power spectrum multipoles measured from the halo catalog of Quijote simulation (see Villaescusa-Navarro et al., 2020, for a description of the simulations). We use the power spectrum multipoles corresponding to $\ell=0,1,2$ in 23 linearly spaced wavenumber ($k$) bins ranging from $0.08\ h^{-1}{\rm Mpc}$ to $2.8\ h^{-1}{\rm Mpc}$ as our observations, $x$. We use the NLE method to learn the likelihood $p(x|\theta)$, where the parameters $\theta$ correspond to three cosmological parameters: the total matter density $\Omega_{m}$, amplitude of the matter fluctuations $\sigma_{8}$, and the Hubble constant $h$ measuring the rate of the expansion of the Universe.
Using the sbi backend, we train a NLE model using an ensemble of six NDEs, each using a thin MAF architecture, with ten hidden layers and three transformations. We use 1800 simulations from Quijote to train the network, retaining the remaining 200 to test its performance as shown in Figure 6. We then use VI sampling to obtain the posterior of the parameters from this trained likelihood ensemble, weighted by their validation losses. This greatly accelerates the sampling process on the full test set when compared to traditional MCMC sampling.
@@ 5.3 Field-level Inference with Graph Neural Networks
We next demonstrate LtU-ILI’s ability to handle complex data modalities by addressing the same cosmological inference problem using a point cloud dataset. Instead of power spectrum summaries, we instead use discrete catalogs of the $10,000$ heaviest dark matter halos in each of the Quijote simulations (Villaescusa-Navarro et al., 2020). The high resolution information present in the discrete halo positions has been shown to constrain cosmological parameters to extremely high precision (e.g., Makinen et al., 2022b; Cuesta-Lazaro and Mishra-Sharma, 2023; de Santi et al., 2023b). We ingest this catalog into our neural architectures using a message-passing graph neural network (Gilmer et al., 2020), which treats dark matter halos as nodes on a graph and filters down field-level information into highly-informative neural representations. This graph information is then fed into flow-based NDEs for performing NPE. We note that usage of graph-type inputs is only possible through the lampe backend in LtU-ILI, as was employed here.
We train this graph-enhanced NPE model to recover the cosmological parameters $\Omega_{m}$ and $\sigma_{8}$ from the positions of Quijote’s heaviest dark matter halos. The graph neural network is a message-passing neural network with node and edge neural networks being two multi-layer perceptrons with $3$ layers of $128$ hidden units each. The network takes as inputs 3D distances between halos, and their corresponding modulus. Nearby halos are connected within the graph if their comoving distance smaller than $20\ h^{-1}{\rm Mpc}$. For NDEs, we train an ensemble of four networks, each one using a MAF architecture with five transformations and $50$ hidden features.
In Figure 7, we show an example point cloud together with the resulting graph neural network inference. The graph neural network obtains informative posteriors, consistent with true values. We note that the constraints are not as tight as those produced for the Quijote dark matter power spectrum in Figure 6. This is likely because the graph only targets the most massive halos, i.e., a more physically realistic probe, while the results in Figure 6 access the full, high-resolution dark matter distribution. In any case, LtU-ILI with the lampe backend provides an easily accessible pathway for new users to integrate exotic architectures such as graph architectures into their development pipeline.
@@ 5.4 Multi-round Inference for Gravitational Waves
(a) Example of GW signal in source and detector frame.
(b) Multi-round training and validation posterior probability.
@@
We next test the performance of LtU-ILI when addressing sequential learning tasks, specifically focusing on gravitational wave reconstruction.. Over recent years, several studies based on ILI methods have addressed the problem of reconstructing parameters of merger events from gravitational wave (GW) signals (Cutler and Flanagan, 1994; Thrane and Talbot, 2019). Initial advancements were marked by significant contributions to posterior estimation using normalizing flows, as highlighted in Green et al. (2020); Green and Gair (2020). Subsequent research (DINGO: Dax et al., 2021a) introduced variations of NPE taking advantage of Group Equivariant Neural Posterior Estimation (GNPE: Dax et al., 2021b) or Flow Matching Posterior Estimation (Wildberger et al., 2023). More recently, Crisostomi et al. (2023) showcased a similar application of multi-round NPE to estimate ringdown parameters of GW150914. Delaunoy et al. (2020) and Bhardwaj et al. (2023) also serve as recent illustrations of applying NRE to GW analyses – the latter incorporating multi-round inference as well.
@@
In this example, we explore the application of multi-round inference to constrain a reduced set of gravitational wave (GW) parameters. Our forward-modeled data vector is constructed by merging the simulated frequency-domain strain signals from each of the Hanford (H1) and Livingstone (L1) detectors. We employ the IMRPhenomPv2 algorithm (Khan et al., 2016) with a frequency resolution of $\Delta f=0.125\textrm{ Hz}$, resulting in an input data vector of $\textrm{dim}(\bm{x})=2\times 7872$. An illustration of this data vector is presented in Figure 8a for clarity. We stress that, in contrast to likelihood-based sampling methods implemented in Bilby (Ashton et al., 2019; Romero-Shaw et al., 2020), which only necessitate the signal component of the forward model for the likelihood computation, our ILI approach requires the inclusion of noise realizations in the training data vector.
@@
We illustrate a simplified application of LtU-ILI for SNPE on GW signals, fixing parameters related to the time and geometry of the detection and focusing on a set of four parameters: chirp mass ($\mathcal{M}$), mass ratio ($q$), effective aligned spin ($\chi_{\textrm{eff}}$) and luminosity distance ($d_{L}$). These parameters, chosen for their reduced correlation and significance in GW parameter estimation (e.g., Veitch et al., 2015; Vitale et al., 2017), are derived from the masses of the coalescing objects – with $\mathcal{M}\equiv(m_{1}m_{2})^{\frac{3}{5}}/(m_{1}+m_{2})^{\frac{1}{5}}$ and $q\equiv m_{2}/m_{1}\leq 1$ – and spin characteristics (Racine, 2008).
@@
We established baseline values for our simulated event $\bm{x_{0}}$ at {$d_{L}=390\ \textrm{Mpc},q=0.79,\mathcal{M}=30.2\ M_{\odot},\chi_{\textrm{eff}}=-0.09$)}, akin to the GW150914 event as detailed in the first column of Table 1 by Abbott et al. (2016). Following the procedure described in Karathanasis et al. (2023), the defined priors are uniform ranging from [100.0 $1.0100.0\textrm{ Mpc},0.2,12.0\,M_{\odot},-$ 1.0] to [1000.0 $1.01000.0\textrm{ Mpc},1.0,45.0\,M_{\odot},$ 1.0], with narrower margins for chirp mass and luminosity distance to facilitate a straightforward example. For density estimation, we utilize the sbi backend with a MAF and a 1D convolutional embedding network, undergoing five rounds of training with 1000 simulations per round.